The mean blood pressure of a large population of adult males is 200 mmHg with a standard deviation of 20 mmHg. Assume that blood pressure measurements are normally distributed. (a) What is the probability that a sample of 100 men from this age group will have mean blood pressure below 204 mmHg? (b) If a group of 25 older men who are strict vegetarians have a mean blood pressure of 188 mmHg, would you say that vegetarianism significantly lowers blood blood levels?
1
Expert's answer
2020-04-20T11:16:45-0400

Based on the central limit theorem, the sample mean is distributed as a normal distribution with mean 200 mmHg and a standard deviation "20/ \sqrt{100}=2" mmHg.


"P(X<204)=P(Z<{204-200\over2})=P(Z<2)\approx0.977250"

"H_0:\mu=200"

"H_1:\mu<200"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


"z={\bar{X}-\mu \over \sigma/\sqrt{n}}={188-200 \over 20/\sqrt{25}}=-3"

The p-value is "p=0.00135" and since "p=0.00135<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population "\mu"  is less than 200, at the 0.05 significance level.

We may say that vegetarianism significantly lowers blood levels.